OpenStudy (anonymous):

Given the linear equation y = -6x – 2, find the y-coordinates of the points (-1, ), (-2, ), and (1, )

OpenStudy (anonymous):

4

OpenStudy (anonymous):

10

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

You understand how to do that, right? All you need to do is plug in those given points for x. E.g. for the first one, -1, y = -6(-1) - 2 = 4.

OpenStudy (anonymous):

Here's how to do these - for example, here the first one. In the point (-1,?) we are told that x=-1 so, we plug (-1) instead of "x" in the equation: y=(-6)*(-1)-2 = +6-2 = 4 y=4 So the answer is: (-1,4)